Isabel Briggs Myers was a pioneer in the study of personality types. The personality types are broadly defined according to four main preferences. Do married couples choose similar or different personality types in their mates? The following data give an indication.

Similarities and Differences in a Random Sample of 375 Married Couples
Number of Similar Preferences Number of Married Couples
All four
Three
Two
One
None
32
133
124
60
26

Suppose that a married couple is selected at random.

(a) Use the data to estimate the probability that they will have 0, 1, 2, 3, or 4 personality preferences in common. (Enter your answers to 2 decimal places.)

0 1    2 3 4
Answer #1

a) The probability table will be:

x Probability
0 26/375 = 0.07
1 60/375 = 0.16
2 124/375 = 0.33
3 133/375 = 0.35
4 32/375 = 0.09